what is the vertex of y=-x2+5x-10
Answer by lwsshak3(11628)

what is the vertex of
y=-x^2+5x-10
complete the square
y=-(x^2-5x+25/4)+25/4-10
y=-(x-5/2)^2-15/4
This is an equation of a parabola that opens down.
Its standard form: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex.
vertex of given parabola:(5/2,-15/4)